# Asymptotic Stability of 3D Out-flowing Compressible Viscous Fluid under Non-Spherical Perturbation

### v1

@@ Appendix A Hardy’s Inequality
Substituting the above into (A), it follows that
@@
By density argument, (A.54) holds for any $u\in H^{1}(\Omega)$. We remark that the trace of $u$ on the boundary domain $\{|x|=1\}$ is well defined by the trace embedding theorem, $\textrm{tr}\vcentcolon H^{1}(\Omega)\to L^{2}(\partial\Omega)$. Taking $R\to\infty$ on the resultant inequality, we obtain that
@@
This concludes the proof. $\square$
Appendix B Spherical Differential Operators
In this appendix, we study the properties of differential operators in spherical coordinate. First, we give the detailed construction of cut-off functions claimed in (3.30). Define the function $\tilde{\xi}(\theta)\in\mathcal{C}^{1}[0,\pi]$ by
$\tilde{\xi}(\theta)\vcentcolon=\left\{\begin{aligned} &0&&\text{if }\ \theta\in[0,\tfrac{\pi}{8})\\ &\tfrac{32}{\pi^{2}}(\theta-\tfrac{\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{\pi}{8},\tfrac{\pi}{4}),\\ &1-\tfrac{32}{\pi^{2}}(\theta-\tfrac{3\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{\pi}{4},\tfrac{3\pi}{8}),\\ &1&&\text{if }\ \theta\in[\tfrac{3\pi}{8},\tfrac{5\pi}{8}),\end{aligned}\right.\quad\left\{\begin{aligned} &1&&\text{if }\ \theta\in[\tfrac{3\pi}{8},\tfrac{5\pi}{8}),\\ &1-\tfrac{32}{\pi^{2}}(\theta-\tfrac{5\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{5\pi}{8},\tfrac{3\pi}{4}),\\ &\tfrac{32}{\pi^{2}}(\theta-\tfrac{7\pi}{8})^{2}&&\text{if }\ \theta\in[\tfrac{3\pi}{4},\tfrac{7\pi}{8}),\\ &0&&\text{if }\ \theta\in[\tfrac{7\pi}{8},\pi].\end{aligned}\right.$
It can be verified that for all $\theta\in[0,\pi]$,
$0\leq\tilde{\xi}(\theta)\leq 1,\qquad|\tilde{\xi}^{\prime}(\theta)|^{2}\leq\frac{128}{\pi^{2}}\xi(\theta),\qquad\text{supp}\big{(}\tilde{\xi}\big{)}=\big{[}\frac{\pi}{8},\frac{7\pi}{8}\big{]}.$ (B.55)
Using the above properties, one can choose a suitable mollifier $\eta\in\mathcal{C}^{\infty}_{c}[-1,1]$ such that the convolution $\xi\vcentcolon=\tilde{\xi}\ast\eta$ satisfies,
$0\leq\xi\leq 1,\qquad|\xi^{\prime}|^{2}\leq C\xi,\qquad\text{supp}(\xi)\subseteq\big{[}\frac{\pi}{9},\frac{8\pi}{9}\big{]},\qquad\sup\limits_{\theta\in[0,\pi]}\big{|}\xi^{(k)}(\theta)\big{|}\leq C,$
for all $k=1,2,3,4$, where $C>0$ is some generic constant. Using this, we define the partition of unity functions as
$\chi_{V}(x)\vcentcolon=\xi\Big{(}\arccos\big{(}\frac{x_{3}}{|x|}\big{)}\Big{)}\quad\text{ and }\quad\chi_{H}(x)\vcentcolon=\xi\Big{(}\arccos\big{(}\frac{x_{2}}{|x|}\big{)}\Big{)}.$
Since $\text{supp}(\xi)=[\frac{\pi}{9},\frac{8\pi}{9}]$, one sees that $\chi_{V}=0$ for $\arccos(\frac{x_{3}}{|x|})\in[0,\frac{\pi}{9}]\cup[\frac{8\pi}{9},\pi]$. The gradient of $\chi_{V}$ is given by:
@@ Appendix B Spherical Differential Operators
Thus using (B.55), we also have for all $x\in\Omega$,
@@
The statements for $\chi_{H}(x)$ is shown following the exact same argument. This proves the existence of functions described in (3.30). In what follows, we use for simplicity, the notations $(\theta,\varphi)\equiv(\theta_{V},\varphi_{V})$ or $(\theta_{H},\varphi_{H})$ and $\chi\equiv\chi_{V}$ or $\chi_{H}$.
Let $(r,\theta,\varphi)\in(0,\infty)\times[0,\pi]\times[0,2\pi)$ be the spherical coordinate. Then the following differential relations for unit basis vectors holds:
$\partial_{r}\hat{r}=0,\qquad$ $\partial_{\theta}\hat{r}=\hat{\theta},\qquad$ $\partial_{\varphi}\hat{r}=\hat{\varphi}\sin\theta,$ (B.56a)
$\partial_{r}\hat{\theta}=0,$ $\partial_{\theta}\hat{\theta}=-\hat{r},$ $\partial_{\varphi}\hat{\theta}=\hat{\varphi}\cos\theta,$ (B.56b)
$\partial_{r}\hat{\varphi}=0,$ $\partial_{\theta}\hat{\varphi}=0,$ $\partial_{\varphi}\hat{\varphi}=-\hat{r}\sin\theta-\hat{\theta}\cos\theta.$ (B.56c)
Let $F$ be a scalar-valued function and $V$ be a vector-valued function. Then the gradient, divergence, and Laplacian under spherical coordinate are given by
$\nabla F=\hat{r}\partial_{r}F+\hat{\theta}\frac{1}{r}\partial_{\theta}F+\hat{\varphi}\frac{1}{r\sin\theta}\partial_{\varphi}F,$ (B.57a)
$\textnormal{{div}}V=\frac{1}{r^{2}}\partial_{r}\big{(}r^{2}(\hat{r}\cdot V)\big{)}+\frac{1}{r\sin\theta}\partial_{\theta}\big{(}(\hat{\theta}\cdot V)\sin\theta\big{)}+\frac{1}{r\sin\theta}\partial_{\varphi}\big{(}\hat{\varphi}\cdot V\big{)},$ (B.57b)
$\Delta F=\frac{1}{r^{2}}\partial_{r}\big{(}r^{2}\partial_{r}F\big{)}+\frac{1}{r^{2}\sin\theta}\partial_{\theta}\big{(}\partial_{\theta}F\sin\theta\big{)}+\frac{1}{r^{2}\sin^{2}\theta}\partial_{\varphi}^{2}F.$ (B.57c)
Proposition B.1.
Let $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$ and $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$. For given $N\in\mathbb{N}$, there exists a constant $C=C(N)>0$ such that for all scalar functions $F(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$,
$\chi\big{|}[\partial^{N},\nabla]F\big{|}\leq C\chi\sum_{m=0}^{N-1}|\nabla\partial^{m}F|,\qquad\chi\big{|}[\partial^{N},\Delta]F\big{|}\leq C\chi\frac{1}{r^{2}}\sum_{m=1}^{N+1}|\partial^{m}F|.$
Moreover, for all $F(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$ and $V(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$,
@@
By product rule, binomial formula, and (B.57a), it holds that
$[\partial^{N},\nabla]F=\sum_{m=0}^{N-1}\binom{N}{m}\Big{\{}\big{(}\partial^{N-m}\hat{r}\big{)}\partial_{r}\partial^{m}F+\frac{\partial^{N-m}\hat{\theta}}{r}\partial_{\theta}\partial^{m}F+\partial^{N-m}\Big{(}\frac{\hat{\varphi}}{r\sin\theta}\Big{)}\partial_{\varphi}\partial^{m}F\Big{\}}.$
Thus using the differential relations (B.56), we obtain the first inequality. Next, by a similar calculation, we also obtain from (B.57c) that
@@
This yields the second inequality of the proposition. Let $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$, and set $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ be a vector-valued function. We have by binomial formula and (B.57)
@@
Multiplying both sides by $\chi$, then taking absolute value, we get
@@
Computing the commutator $[\partial^{k},\nabla\textnormal{{div}}]$ using the above expression, we obtain the last inequality of the proposition. $\square$
Proposition B.2.
Let $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$ and $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$. Then for $k,l\in\mathbb{N}\cup\{0\}$, there exists a constant $C=C(k,l)>0$ such that, for all $V(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$,
@@
Using differential relations (B.56) and (B.57), we get
@@
Taking the difference of the above two equations, we see that the term $\hat{r}\cdot\partial_{r}^{2}V$ cancels. Then, multiplying $\chi$ on both side of the resultant equality, integrating in $x\in\Omega$, and applying Proposition A.1 on terms of the form $\int_{\Omega}\frac{|V|^{2}}{r^{2}}{\mathrm{d}}x$, we obtain the desired inequality. $\square$

### v2

@@ A.1 Hardy’s inequality
Substituting the above inequality in (A.1), it follows that
@@
By density argument, (A.2) holds for any $u\in H^{1}(\Omega)$. We remark that the trace of $u$ on the boundary domain $\{|x|=1\}$ is well defined by the trace embedding theorem. Taking $R\to\infty$ on the resultant inequality, we obtain that
@@
This completes the proof. $\square$
A.2 Spherical differential operators
In this appendix, we summarize the properties of differential operators in spherical coordinate. First, we give the detailed construction of the cut-off functions claimed in (2.23). Define the function $\tilde{\xi}(\theta)\in\mathcal{C}^{1}[0,\pi]$ by
[…]
We verify that for $\theta\in[0,\pi]$,
$0\leq\tilde{\xi}(\theta)\leq 1,\qquad|\tilde{\xi}^{\prime}(\theta)|^{2}\leq\frac{128}{\pi^{2}}\xi(\theta),\qquad\text{supp}\big(\tilde{\xi}\big)=\big[\frac{\pi}{8},\frac{7\pi}{8}\big].$
Using the above properties, we choose a suitable mollifier $\eta\in\mathcal{C}^{\infty}_{c}[-1,1]$ such that the convolution $\xi\vcentcolon=\tilde{\xi}\ast\eta$ satisfies,
$0\leq\xi\leq 1,\qquad|\xi^{\prime}|^{2}\leq C\xi,\qquad\text{supp}(\xi)\subseteq\big[\frac{\pi}{9},\frac{8\pi}{9}\big],\qquad\sup\limits_{\theta\in[0,\pi]}\big|\xi^{(k)}(\theta)\big|\leq C,$ (A.3)
for $k=1,2,3,4$, where $C>0$ is a generic constant. Using this, we define the partition of unity functions as
$\chi_{V}(x)\vcentcolon=\xi\Big(\arccos\big(\frac{x_{3}}{|x|}\big)\Big)\qquad\text{ and }\qquad\chi_{H}(x)\vcentcolon=\xi\Big(\arccos\big(\frac{x_{2}}{|x|}\big)\Big).$
Since $\text{supp}(\xi)=[\frac{\pi}{9},\frac{8\pi}{9}]$, we see that $\chi_{V}=0$ for $\arccos(\frac{x_{3}}{|x|})\in[0,\frac{\pi}{9}]\cup[\frac{8\pi}{9},\pi]$. The gradient of $\chi_{V}$ is given by
@@ A.2 Spherical differential operators
Thus using (A.3), we also have for $x\in\Omega$,
@@
The assertions for $\chi_{H}(x)$ is shown following the exact same argument above. This proves the existence of functions described in (2.23).
In what follows, we use the abbreviated notations $(\theta,\varphi)\vcentcolon=(\theta_{V},\varphi_{V})$ or $(\theta_{H},\varphi_{H})$ and $\chi\vcentcolon=\chi_{V}$ or $\chi_{H}$. Let $(r,\theta,\varphi)\in(0,\infty)\times[0,\pi]\times[0,2\pi)$ be the spherical coordinate. Then the following differential relations for unit basis vectors holds:
[…]
$\partial_{r}\hat{\theta}=0,$ $\partial_{\theta}\hat{\theta}=-\hat{r}$, $\partial_{\varphi}\hat{\theta}=\hat{\varphi}\cos\theta,$ (A.4b)
$\partial_{r}\hat{\varphi}=0,$ $\partial_{\theta}\hat{\varphi}=0,$ $\partial_{\varphi}\hat{\varphi}=-\hat{r}\sin\theta-\hat{\theta}\cos\theta.$ (A.4c)
Let $F\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$ be a scalar-valued function and $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ be a vector-valued function. Then the gradient, divergence, and Laplacian under spherical coordinate are given by
$\nabla F=\hat{r}\partial_{r}F+\hat{\theta}\frac{1}{r}\partial_{\theta}F+\hat{\varphi}\frac{1}{r\sin\theta}\partial_{\varphi}F,$ (A.5a)
$\textnormal{{div}}V=\frac{1}{r^{2}}\partial_{r}\big(r^{2}(\hat{r}\cdot V)\big)+\frac{1}{r\sin\theta}\partial_{\theta}\big((\hat{\theta}\cdot V)\sin\theta\big)+\frac{1}{r\sin\theta}\partial_{\varphi}\big(\hat{\varphi}\cdot V\big),$ (A.5b)
$\Delta F=\frac{1}{r^{2}}\partial_{r}\big(r^{2}\partial_{r}F\big)+\frac{1}{r^{2}\sin\theta}\partial_{\theta}\big(\partial_{\theta}F\sin\theta\big)+\frac{1}{r^{2}\sin^{2}\theta}\partial_{\varphi}^{2}F.$ (A.5c)
Proposition A.2.
Let $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$ and $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$. For an arbitrary non-negative integer $N$, there exists a constant $C=C(N)>0$ such that for an arbitrary scalar functions $F(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$, we have
$\chi\big|[\partial^{N},\nabla]F\big|\leq C\chi\sum_{m=0}^{N-1}|\nabla\partial^{m}F|,\hskip 18.49988pt\chi\big|[\partial^{N},\Delta]F\big|\leq C\chi\frac{1}{r^{2}}\sum_{m=1}^{N+1}|\partial^{m}F|.$
Moreover, for functions $F(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}$ and $V(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$,
@@
By product rule, the binomial theorem and (A.5a), it holds that
$\partial^{N},\nabla]F=\sum_{m=0}^{N-1}\binom{N}{m}\Big\{\big(\partial^{N-m}\hat{r}\big)\partial_{r}\partial^{m}F+\frac{\partial^{N-m}\hat{\theta}}{r}\partial_{\theta}\partial^{m}F+\partial^{N-m}\Big(\frac{\hat{\varphi}}{r\sin\theta}\Big)\partial_{\varphi}\partial^{m}F\Big\}.$
Thus using the differential relations (A.4), we obtain the first inequality. By a similar calculation, we also obtain from (A.5c) that
@@
This yields the second inequality of the proposition. Let $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$ and $V\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$ be a vector-valued function. We have by the binomial theorem and (A.5) that
@@
Multiplying both sides by $\chi$ and taking the absolute value, we get
@@
Computing the commutator $\partial^{k},\nabla\textnormal{{div}}]$ using the above expression, we obtain the last inequality in the proposition. $\square$
Proposition A.3.
Let $\partial\in\{\partial_{\theta},\partial_{\varphi}\}$ and $D\in\{\partial_{r},\partial_{\theta},\partial_{\varphi}\}$. Then for $k,l\in\mathbb{N}\cup\{0\}$, there exists a constant $C=C(k,l)>0$ such that for an arbitrary vector-valued function $V(x)\vcentcolon\mathbb{R}^{3}\to\mathbb{R}^{3}$, it holds that
@@
Using differential relations (A.4) and (A.5), we get
@@
Taking the difference between the above two equations, we see that the term $\hat{r}\cdot\partial_{r}^{2}V$ cancels. Using this and Proposition A.1, we obtain the desired estimate. $\square$